Find the perimeter of a right triangle with legs measuring 7 and 24.
Answer by benni1013(206)

A couple of pretexts:
1) Perimeter deals with the measure around the shape NOT inside the shape
2) You need to use the pythagorean theorem (a^2+b^2=c^2). A and B are the legs. C is the hypotenuse, the long side.
First, what do we know.
1) We know the legs of the triangle a and b equal 7 and 24.
2) Start plugging into the Pythagorean theorem.
WORK:
7^2+24^2=c^2>>49+576=c^2>>625=c^2>>25=c*
*To undo a exponent use a square root and vice versa
With the hypotenuse found, add the three sides to find the perimeter
WORK:
7+24+25=56 ?*
*For future benefit, ALWAYS label your solutions!